### v1

@@ VIII-B1 LLM-enabled planning in telecom
Telecom networks are complicated large-scale systems, and many tasks require multi-step thinking and planning. For instance, the previous Section IV-C2 introduced an example of coding wireless projects with multi-step planning, and many optimization problems with multiple control variables have to been solved step-by-step[2]. However, recent benchmarks have shown that LLM models struggle with tasks requiring complex planning and sequential decision-making. Some existing studies such as [14] and [18] propose to improve the multi-step planning capabilities by step-by-step and CoT prompting. Despite the satisfactory performance in [14] and [18], they require dedicated analyses to manually decompose a complicated task into multiple sub-tasks. Therefore, future studies should aim at developing better algorithms for planning that can be integrated into LLM models, and such multi-step planning capability is crucial for solving telecom-domain tasks. This might involve incorporating structured reasoning and problem-solving frameworks into the models, enabling them to break down tasks into smaller and more manageable sub-tasks. Therefore, automated task decomposition can be an attractive solution to improve the planning performance of LLM models. However, automatically decoupling one complicated task into multiple sub-tasks is still very challenging in the telecom field. Additionally, another solution could be integrating simulation environments directly within the training process, allowing models to practice and refine their planning skills in a controlled setting before applying them to real-world tasks. It allows the LLM models to improve the planning performance by trial-and-error before applying it to telecom tasks.
@@
VIII-B2 Model compression and fast inference for network edge and mobile applications
@@
The model size is one of the key bottlenecks of applying LLMs to the telecom domain, leading to stringent requirements for computational and storage capacities. Therefore, compressing the model size to adapt to network edge and mobile applications becomes a promising direction. In addition, it will also contribute to the fast inference of LLM models, since many wireless applications require rapid response time and low latency. For instance, Xu et al. proposed an on-device inference model specifically designed for efficient generative natural language processing tasks[97], achieving a 9.3×9.3$\times$9.3 × faster generation speed. Such a technique can be very promising for LLM-enabled mobile applications in telecom, enabling faster response time for user inquiries. Meanwhile, it is worth noting that compressing the model size may degrade the LLM performance, and how to balance the model size and performance requires more research efforts. It calls for novel model compression and pruning techniques to reduce the storage and computation burdens at the network edge; on the other hand, standard metrics must be defined to evaluate the performance of LLM models in the telecom domain, e.g., accuracy and hallucination probability as we introduced in previous Section III-E.
VIII-B3 Overcoming hallucination problems in telecom applications
@@
VIII-B4 Retrieval augmented-LLM for telecom
@@
Retrieval augmentation is an important direction for LLM development, which retrieves facts from an external knowledge base to ground LLM models on the most up-to-date information. Telecom networks are constantly evolving and updating, and retrieval augmentation has great potential for telecom applications. In particular, retrieval-augmented LLM can improve the quality and relevance of the generated responses since the LLM has access to more accurate and relevant information. However, current retrieval-augmented generation models increase the context length, which in turn decreases the efficiency of the model due to the added computational cost, which may lead to severe slow-response issues. Such slow response may increase the overall network latency and degrade the service quality. It may prevent the application of some scenarios with tight delay budget, which is very common in wireless networks. Therefore, future research could focus on improving the efficiency of retrieval-augmented generation by optimizing retrieval mechanisms to balance context relevance and length. This could involve developing more advanced indexing and search algorithms that require less memory. Additionally, dynamically adjusting the amount of retrieved information based on the query’s complexity could help maintain or improve efficiency without sacrificing the quality of the output.
VIII-B5 Economic and affordable LLM models
@@
Despite the great potential and advantages, training an LLM model can be financially expensive. For instance, the training expenses for GPT-4 exceeded \$ 100 million, and the LLaMa2 70B model was trained on 2048 GPUs A100 for 23 days with \$ 1.7 million estimated cost [205]. Although training some smaller models such as LLaMa2 7B can be much cheaper, the affordability of LLM techniques is still one of the main concerns. For instance, the study in [206] shows that using GPT-4 to support customer service can cost more than \$ 21,000 per month for a small business. Meanwhile, there are many LLM APIs with various prices, including the prompt cost proportional to the prompt length, generation cost related to the generation length, and a possible fixed cost per query. For example, it costs \$30 for 10M tokens using OpenAI’s GPT-4, while only \$ 0.2 for GPT-J hosted by Textsyth[207]. The financing cost of training, fine-tuning, and deploying LLM models will significantly affect the application in telecom networks. Given the heterogeneous prices and service quality, it is of great importance to evaluate the financing cost of deploying LLM models in telecom networks, e.g., balancing the possible performance improvement and the LLM deployment cost, and using LLM models in an economic manner for telecom applications. However, this direction has limited existing studies, and it still requires more research efforts.
@@ IX Conclusion
Recently, large language model (LLM) have shown great promise in many fields, specifically for language-related tasks such as summarization and question and answering. LLM-based solutions have also been primarily investigated in the telecom field. In this work, we aim to present a comprehensive survey on LLM for Telecom. In particular, we first introduced the LLM fundamentals. We present a comprehensive overview of the model architecture, pre-training, fine-tuning, inference and utilization, evaluation, and deployment of LLM-based solutions. Then, a comprehensive survey of existing works on the key techniques and applications in terms of generation, classification, optimization, and prediction problems is presented. These investigations and analyses have proven that LLMs have outstanding potential to bring artificial general intelligence to the telecom field using in-context and zero-shot learning capabilities. Finally, we discussed the key challenges, such as data sets and cost, as well as future research opportunities of LLM-empowered telecom. We hope this work can serve as a good reference for researchers and engineers to better understand the existing works, potentials, challenges, and opportunities of applying LLM for the telecom field.

### v2

@@ VIII-B2 LLM-enabled planning in telecom
Multi-step planning and scheduling are critical for handling many tasks in the telecom field. For instance, Section IV-C2 has introduced an example of coding wireless projects with step-by-step prompting. Meanwhile, many optimization problems with multiple network elements and control variables have to be solved by dedicated planning [7]. However, recent benchmarks have shown that LLMs struggle with tasks requiring complex planning and sequential decision-making, which may prevent the direct application to many telecom tasks. Some existing studies such as [14] and [18] propose to improve the multi-step planning capabilities by step-by-step and CoT prompting. Despite the satisfactory performance in [14] and [18], they require dedicated analyses to manually decompose a complicated task into multiple sub-tasks. Therefore, future studies should aim at developing better algorithms for planning that can be integrated into LLMs, and such multi-step planning capability is crucial for solving telecom-domain tasks. This might involve incorporating structured reasoning and problem-solving frameworks into the models, enabling them to break down tasks into smaller and more manageable sub-tasks. Therefore, automated task decomposition can be an attractive solution to improve the planning performance of LLMs. However, automatically decoupling one complicated task into multiple sub-tasks is still very challenging in the telecom field. Additionally, another solution could be integrating simulation environments directly within the training process, allowing models to practice and refine their planning skills in a controlled setting before applying them to real-world tasks. It allows the LLMs to improve the planning performance by trial-and-error before applying it to telecom tasks.
VIII-B3 LLM for resource allocation and network optimization
Resource management is a fundamental and crucial problem for network operation, e.g., transmission power and bandwidth resources allocation. The above Section VI introduced various LLM-enabled optimization techniques for telecom applications, including LLM-aided reinforcement learning, black-box optimizer, LLM-aided convex optimization and heuristic algorithm design. These analyses of existing studies have revealed the potential of using LLM to optimize network performance. For instance, verbal reinforcement learning can take the network operator’s human language instructions as input to improve task performance, and LLMs can design novel heuristic algorithms based on specific task demands for network resource allocations. LLM-based optimization has two crucial advantages: Firstly, LLM can integrate human languages into the optimization procedure, which makes network management more accessible with much lower complexity; Secondly, LLM can provide reasons and explanations for their decisions, and this capability is crucial for understanding complicated systems such as telecom networks. Despite the advantages, it is worth noting that some network optimization problems can be extremely complicated with coupled control variables and correlated network elements. Solving these optimization tasks may require dedicated design and multi-step scheduling, which is still a challenge in the LLM field.
VIII-B4 LLM-enhanced machine learning for telecom
ML algorithms have been widely applied to wireless networks and achieve satisfactory performance. For example, reinforcement learning is one of the most widely used ML techniques for network optimization, and deep neural networks have been extensively explored to predict CSI. Section VI-B introduced LLM-enhanced reinforcement learning by automating the reward function design, indicating a promising direction to explore LLM-enhanced ML algorithms. For instance, Sahu et al. investigated LLM-aided semi-supervised learning for the task of extractive text summarization, in which they proposed a prompt-based pseudo-labelling strategy with LLMs[224]. Multi-agent learning also has many crucial applications in wireless networks, and existing studies show that LLM-based multi-agents have many promising features [225]. In summary, LLM brings new opportunities to make conventional ML algorithms more accessible and explainable when applied to telecom networks.
VIII-B5 Real-world implementations of LLM for telecom industry
The study in [106] by Apple introduced a method for efficiently running LLMs on devices with limited DRAM capacity. This advancement is particularly beneficial for telecom-specific applications that rely on on-device LLM. By bringing LLM capabilities onto the device, Qualcomm also integrates on-device AI into smartphones to provide faster and more personalized services without relying on cloud-based solutions[107]. Qualcomm aims to improve user privacy and reduce latency, enabling applications such as real-time language translation, advanced camera features, and direct contextual assistance on the user’s device. On-device AI also allows for continuous operation without internet dependency, which is crucial for maintaining service quality in areas with poor connectivity. Additionally, solutions such as Kinetica’s SQL-GPT enable telecom professionals to interact with data using natural language queries, converting these queries into SQL for quick and effective analysis[226]. This approach democratizes access to data insights, empowering employees to make faster and more informed decisions. These applications demonstrate the transformative potential of LLMs in the telecom industry, enhancing security, operational efficiency, and customer experience. By continuing to innovate with LLMs, telecom companies can stay ahead in an increasingly AI-driven landscape, providing superior services and maintaining a competitive advantage.
@@ VIII-B5 Real-world implementations of LLM for telecom industry
VIII-B6 Model compression and fast inference for network edge and mobile applications
@@
The model size is one of the key bottlenecks of applying LLMs to the telecom domain, leading to stringent requirements for computational and storage capacities. Therefore, compressing the model size to adapt to network edge and mobile applications becomes a promising direction. In addition, it will also contribute to the fast inference of LLMs, since many wireless applications require rapid response time and low latency. For instance, Xu et al. proposed an on-device inference model specifically designed for efficient generative natural language processing tasks[112], achieving a 9.3×9.3$\times$9.3 × faster generation speed. Such a technique can be very promising for LLM-enabled mobile applications in telecom, enabling faster response time for user inquiries. Meanwhile, it is worth noting that compressing the model size may degrade the LLM performance, and how to balance the model size and performance requires more research efforts. It calls for novel model compression and pruning techniques to reduce the storage and computation burdens at the network edge; on the other hand, standard metrics must be defined to evaluate the performance of LLMs in the telecom domain, e.g., accuracy and hallucination probability as we introduced in previous Section III-E.
VIII-B7 Overcoming hallucination problems in telecom applications
@@
VIII-B8 Retrieval augmented-LLM for telecom
@@
Retrieval augmentation is an important direction for LLM development, which retrieves facts from an external knowledge base to ground LLMs on the most up-to-date information. Telecom networks are constantly evolving and updating, and retrieval augmentation has great potential for telecom applications. In particular, retrieval-augmented LLM can improve the quality and relevance of the generated responses since the LLM has access to more accurate and relevant information. However, current retrieval-augmented generation models increase the context length, which in turn decreases the efficiency of the model due to the added computational cost, which may lead to severe slow-response issues. Such slow response may increase the overall network latency and degrade the service quality. It may prevent the application of some scenarios with tight delay budget, which is very common in wireless networks. Therefore, future research could focus on improving the efficiency of retrieval-augmented generation by optimizing retrieval mechanisms to balance context relevance and length. This could involve developing more advanced indexing and search algorithms that require less memory. Additionally, dynamically adjusting the amount of retrieved information based on the query’s complexity could help maintain or improve efficiency without sacrificing the quality of the output.
VIII-B9 Economic and affordable LLMs
@@
Despite the great potential and advantages, training an LLM can be financially expensive. For instance, the training expenses for GPT-4 exceeded \$ 100 million, and the LLaMa2 70B model was trained on 2048 GPUs A100 for 23 days with \$ 1.7 million estimated cost [228]. Although training some smaller models such as LLaMa2 7B can be much cheaper, the affordability of LLM techniques is still one of the main concerns. For instance, the study in [229] shows that using GPT-4 to support customer service can cost more than \$ 21,000 per month for a small business. Meanwhile, there are many LLM APIs with various prices, including the prompt cost proportional to the prompt length, generation cost related to the generation length, and a possible fixed cost per query. For example, it costs \$30 for 10M tokens using OpenAI’s GPT-4, while only \$ 0.2 for GPT-J hosted by Textsyth[230]. The financing cost of training, fine-tuning, and deploying LLMs will significantly affect the application in telecom networks. However, advancements like OpenAI’s GPT-4o mini, a cost-efficient small model, offer promising solutions. Priced at just \$ 0.15 per million input tokens and \$ 0.60 per million output tokens, GPT-4o mini is more affordable than previous frontier models and over 60% cheaper than GPT-3.5 Turbo. This affordability, combined with its superior performance in reasoning tasks, mathematical reasoning, and coding proficiency, enables a broad range of cost-effective applications. Similarly, Llama3-8b is also an affordable small-scale model with fast inference speed. These small-scale models may alleviate the economic cost of LLMs with fewer parameters, lower training and fine-tuning costs, and faster inference time. For instance, telecom companies can use GPT-4o mini for customer support chatbots that handle vast conversation histories or for network management systems that analyze extensive performance metrics in real time. Given the heterogeneous prices and service quality, it is of great importance to evaluate the financing cost of deploying LLMs in telecom networks, e.g., balancing the possible performance improvement and the LLM deployment cost, and using LLMs in an economic manner for telecom applications. However, this direction has limited existing studies, and it still requires more research efforts.
@@ IX Conclusion
Recently, large language models (LLMs) have shown great promise in many fields, specifically for language-related tasks such as summarization and question and answering. LLM-based solutions have also been primarily investigated in the telecom field. In this work, we aim to present a comprehensive survey on LLMs for Telecom. In particular, we first introduced the LLM fundamentals. We present a comprehensive overview of the model architecture, pre-training, fine-tuning, inference and utilization, evaluation, and deployment of LLM-based solutions. Then, a comprehensive survey of existing works on the key techniques and applications in terms of generation, classification, optimization, and prediction problems is presented. These investigations and analyses have proven that LLMs have outstanding potential to bring artificial general intelligence to the telecom field using in-context and zero-shot learning capabilities. Finally, we discussed the key challenges, such as data sets and cost, as well as future research opportunities of LLM-empowered telecom. We hope this work can serve as a good reference for researchers and engineers to better understand the existing works, potentials, challenges, and opportunities of applying LLMs for the telecom field.